Item to Classify Standard Actual Type of Variance
Sales volume 42,700 uni ts 44, 600 uni
ts F
Sa
les price $3.70 per unit $3.67 per unit
Materials cost $3.70 per pound $3.60 per pound
Materials usage 90,700 pounds 91,700 pounds
Labor cost $10.30 per hour $10.70 per hour
Labor usage 62,500 hours 61,700 hours
Fixed cost spending $495,320 $459,380
Fixed cost per unit (volume) $11.60 per unit $10.30 per unit

Indicate whether each of the following variances is favorable (F) or unfavorable (U). The first one has been done as an example. (Select "None" if there is no effect (i.e., zero variance).)
Answer #1
Item to Classify Standard Actual Variance ( Standard - Actual) Type of Variance
Sales volume 42,700 units 44,600 units (42,700-44,600) = $1,900 units Favorable
Sales price $3.70 per unit $3.67 per unit (3.70-3.67) = $0.03 per unit Unfavorable
Materials cost $3.70 per pound $3.60 per pound (3.70-3.60) = $0.10 per pound Favorable
Materials usage 90,700 pounds 91,700 pounds (90,700-91,700) = $1,000 pounds Unfavorable
Labor cost $10.30 per hour $10.70 per hour (10.30-10.70) = $0.40 per hour Unfavorable
Labor usage 62,500 hours 61,700 hours (62,500-61,700) = 800 hours Favorable
Fixed cost spending $495,320 $459,380 (495,320-459,380) = $35,940 Favorable
Fixed cost per unit (volume) $11.60 per unit $10.30 per unit (11.60-10.30) = $1.30 per unit Favorable

Kindly comment if you need further assistance. Thanks‼!